Learning Object-Oriented Technology. When and How ?

Maximo Prieto

Departmento de Informatica Universidad Nacional de La Plata Buenos Aires, Argentina

Abstract

In this lecture we will focus on several aspects of the learning process of Object-Oriented concepts. We will discuss issues such as: Which is the best moment (in the c.s. curricula) to begin learning about objects. Which are the best tools for supporting this process.

In particular we will talk about our experience in 10 universities in Argentina, Uruguay, Chile and Ecuador, in different stages of the career (undergraduate, graduate).

We will talk about current projects on Object-Orientation at LIFIA, Argentina: building animations of Object-Oriented architectures and using hypermedia in software engineering environments.